INAF queue observing, 2018 Jan 29-30

We (the observers) take this space to bring attention of the world to the the awkward situation of large part of research personnel in INAF. We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world. But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.

However, we believe that like for many other INAF post-docs, our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing. Like many other post-docs, we are paid on contracts renewed every year pending the availability of external funds.

Moreover, we are not entitled to officially say a word on institutional choice and in many aspects we are not considered as INAF employees, but like external collaborators. Indeed, we cannot even lead INAF projects and build a independent research.

Now, we believe INAF has given the power and instruments to start changing this situation.

Observer: R. Carini, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: B.
Instrument: LUCI2

Summary

At the sunset the sky is clear.

instruments: LBC works, LUCI1 only imaging and AO, LUCI2 imaging and LS and MOS, MODS1 works, MODS2 only blue channel

Night duration (12\xB0 el twilight) 2:00am- 7:30pm UT = 5 h 30m THIS IS A SHARED NIGHT

Observing time: 4 h 45m (86%)

Weather time loss: * *h m (%)

technical time loss 0.75h m (14%) ()

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Darks Antoniucci LUCI2      
Vanzella JDz11 LUCI1-AO Ks 2h 15m field A, 2m field B  

Details

*luci dark dit 200, ndit 1, mer, norm luci2.20180130.0005-0006 * dit 300 ndit 1, mer, norm luci2.20180130.0008-0011

02:00 begin with AO, Vanzella program
  • 02:07 sent the preset
  • 02:10 preset failed, the TO configures again the rotator
  • 02:13 sent the preset
  • luci1.20180130.00146 no good, something failed in AO
  • 02:21we send again the preset, but it failed
  • 02:22 we send the preset again, seeing about 1.00''on the guider and DIMM
  • measured mag on the ref star 13.3, freq 400, binning 2.
  • ==luci1.0147-0151==
  • luci1.0147 fwhm = 0.24'' (upper limit, measured on a galaxy)
  • =luci1.0150 fwhm = 0.23'' (upper limit)
  • 02:36 error

02:39 we can continue the integration
  • ==luci1.0152-0166==
  • luci1.0157 fwhm 0.19'' (upper limit)
  • luci.0160 fwhm 0.23 , seeing about 0.9''
  • pass to field B with the PSF star luci1.0167-0169
  • luci1.0167 fwhm = 0.17''
  • 03:05 error
  • 03:07 resume but it doesn't work

03:08 sent the script again, failed
  • 03:09 we send again the preset
  • ==luci1.0170-0189==
  • luci1.0170 fwhm= 0.2''(upper limit), seeing about 1.0''
  • luci1.0175 fwhm = 0.22'' (upper limit)
  • 03:29, error in the AO, we put pause
  • 03:30 restart the integration
  • luci1.0180 fwhm =2.18'' (upper limit), seeing about 0.9''
  • 03:30 loop fault, guide lost
  • restart the integration
  • luci1.0184 fwhm = 0.23''
  • filed B with the PSF star luci1.0190-0193
  • luci1.0192 fwhm =0.19''

03:57 sent the script again, but the preset failed
  • we send again the preset, but failed
  • 04:00 the guide star it is almost out from the field of view, TO checks pointing

04:02 start again the script
  • ==luci1.0194-0213==
  • image 197 not good lost AO for few seconds
  • luci1.0198 fwhm =0.2''(upper limit) seeing 0.9''
  • luci1.0204 fwhm=0.22 seeing about 1.0 on the DIMM
  • luci.0206 fwhm=0.17''
  • pass to field B with the PSF star luci1.0214-0219
  • 04:39 offset failed
  • 04:40 continue
  • luci1.0214 fwhm 0.2'', seeing 0.8 on the guider
  • 04:45 script finished for the first time!!!!!!

04:45 we send again the script
  • ==luci1.0220-0225==
  • luci1.0224 fwhm=0.18'''
  • 04:55 lost guide star

04:57 sent the preset again ==luci1.0226-==
  • 05:03 loop closed start the integration
  • 05:04 loop suspended, == the image 0227 could be not good
  • 05:05 start again
  • 05:11 AO failed

05:15 sent again the preset
  • 05:20 start the integration ==luci1.0228-0230==
  • 05:21 failed
  • 05:22 start again the integration ==luci1.0231-0242==
  • luci1.0233 fwhm= 0.2'', seeing 0.9''on the guider
  • 05:30 loop fault
  • 05:31 start again
  • luci1.0239 fwhm=0.2''seeing about 0.7''-0.8'''
  • 05:35 loop failed
  • 05:50 send preset again
  • 05:57 sent the preset again
  • 05:58 and again
  • 06:02 start the integration ==luci1.0243-0262== seeing stable less than 1.0'' on the guider , 25 MIN TEC LOSS
  • luci1.0246 fwhm= 0.21''
  • luci1.0258 fwhm= 0.18''
  • 06:33 field B luci1.0263-0268
  • luci1.0263 fwhm = 0.19''
  • we have finished again the script!!yuppiiiiiii!!!!!!

06:37 sent again the script ==luci1.0269-0287==
  • 06:41 the seeing on the guider is about 1''-1.2''
  • luci1.0274 fwhm =0.26''
  • 06:55 the seeing is good again, about 0.8''-0.9''
  • field B luci1.0288-0293
  • luci1.028 fwhm= 0.22'''
  • 07:07 finish again the script (third!!!!!!!)wow!!!!!

07:08 sent the script again
  • ==luci1.0294-=0307=
  • luci1.0300 fwhm=0.25''
  • field B luci1.0308-0310
  • luci1.0310 fwhm =0.15'''

-- %USERSIG{AndreaRossi - 2018-01-30}%

Comments

 
Topic revision: r5 - 14 Feb 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback